A galvanometer with a resistance of 15.0Ω gives a full-scale reading for current of 3.0 A. What shunt resistance is needed to convert this into an ammeter with a 30.0 A full-scale reading?

1
Expert's answer
2022-02-18T08:23:17-0500

According to Ohm's law, the currents in parallel circuits are inversely proportional to the resistances:


"\frac{I_a}{I_{sh}}=\frac{R_{sh}}{R_a}",


where "I_a" - the current through the ammeter,

"I_{sh}" - the current through the shunt resistance,

"R_a" - the resistance of the ammeter,

"R_{sh}" - the shunt resistance.


The current through the shunt resistance:


"I_{sh}=I_{fs}-I_a" ,


where "I_{fs}" - the current of full-scale reading.


So,


"R_{sh}=R_a\frac{I_a}{I_{fs}-I_a}=15\frac{3}{30-3}=1.67\space Ohm".